### v1

# Optimizing Negative Prompts for Enhanced Aesthetics and Fidelity in Text-To-Image Generation

Abstract
In text-to-image generation, using negative prompts, which describe undesirable image characteristics, can significantly boost image quality. However, producing good negative prompts is manual and tedious. To address this, we propose NegOpt, a novel method for optimizing negative prompt generation toward enhanced image generation, using supervised fine-tuning and reinforcement learning. Our combined approach results in a substantial increase of 25% in Inception Score compared to other approaches and surpasses ground-truth negative prompts from the test set. Furthermore, with NegOpt we can preferentially optimize the metrics most important to us. Finally, we construct Negative Prompts DB, a dataset of negative prompts.
1 Introduction
Prompting has become the standard approach for adapting large models to specific tasks Liu et al. (2023), with prompt optimization techniques Lester et al. (2021); Deng et al. (2022); Li and Liang (2021); Liu et al. (2022) playing a crucial role in enhancing performance. This trend is also evident in text-to-image generation Ramesh et al. (2022). Despite recent improvements in image-generation quality Rombach et al. (2022), challenges remain. In this work, we look to enhance image-generation aesthetics and fidelity by optimizing negative prompts.
Images produced by text-to-image generation models Ramesh et al. (2022); Rombach et al. (2022) sometimes suffer from issues such as blurriness and poor framing. Negative prompts can divert models from such unwanted characteristics. Wong (2023) shows that negative prompts with phrases like “blurriness” and “out of frame”, can be more effective than normal prompts with phrases such as “clear” and “good framing” in preventing unwanted characteristics (see Figure 1).
| NegOpt | Baseline | NegOpt | Baseline |
Recent work has attempted to improve image quality by optimizing normal prompts Hao et al. (2022). Prior to this, however, optimizing prompts for image generation was performed manually in a tedious and error-prone process.
To address these issues, we propose a novel approach called NegOpt, which performs negative prompt optimization through a two-step process. First, we treat the task of generating prompts as a sequence-to-sequence (seq2seq) language modeling problem, where the input is a normal prompt, and the output is a negative prompt, . We fine-tune a seq2seq model on a first-of-its-kind dataset of - pairs, which we construct and call Negative Prompts DB. Second, we frame the task as a reinforcement learning problem, maximizing a scalar reward signal based on the quality of images generated with and .
Our results demonstrate a significant improvement in generated-image aesthetics and fidelity with a 25% increase in Inception Scores as well as improvements in other metrics. We even surpass the ground-truth negative prompts when evaluating on the test set. Moreover, we show that NegOpt improves our most valued metrics, which are aesthetics and fidelity, without sacrificing others such as prompt alignment. Our key contributions are:
1.
We introduce NegOpt, a method for optimizing negative prompts for image generation that improves aesthetics and fidelity.
2.
We construct Negative Prompts DB, a dataset specifically aggregating negative prompts.
2 Background
Recent research has made significant progress in optimizing prompts for various tasks, mainly focusing on gradient-based methods Qin and Eisner (2021); Li and Liang (2021); Liu et al. (2022) and reinforcement learning approaches Deng et al. (2022). However, most of this work targets pure natural language processing tasks.
In text-to-image generation, one study has successfully explored optimizing normal prompts Hao et al. (2022), but negative prompt optimization remains unaddressed. This gap is particularly noteworthy considering the extensive use of negative prompts in online communities111https://stable-diffusion-art.com/how-to-use-negative-prompts/.
Prompt |
Negative Prompt |
Image |
Other Images |
A Wolf, high detailed fur, monochrome charcoal painting art, detail work, various tones of black, blended, colorful detailed eyes, by Gerd Amble, 8k, 2d render |
None |
Ground Truth |
multiple heads, multiple wolfs, close up |
NegOpt |
elongated body, cropped image, red eyes, out of frame, draft, deformed hands, signatures, big hair, twisted fingers, double image, long neck, malformed hands, multiple heads, extra limb, ugly… |
3 NegOpt
Our contributions consist of a unique dataset and a novel method for optimizing negative prompts.
3.1 Dataset: Negative Prompts DB
We compiled a collection of publicly-available prompt samples from Playground AI222https://playgroundai.com/, selecting posts submitted between December 1, 2022, and March 7, 2023, that received at least one like. This resulted in a total of 256,224 samples, collected on March 9, 2023. The following attributes are available: id, created_at, model, sampler, likes, height, steps, width, cursor, url, and cfg_scale. In our experimental setup, we describe how we use subsets of Negative Prompts DB within our experiments (see Section 4.1 and Section 4.2).
3.2 Core Method: NegOpt
NegOpt consists of two phases: Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL) toward producing optimized negative prompts.
3.2.1 Phase 1: Supervised Fine-tuning (SFT)
In this phase, we fine-tune a sequence-to-sequence (seq2seq) model Sutskever et al. (2014) using a subset of the Negative Prompts DB dataset (see Section 4.1). We set the normal prompt as the source, and the negative prompt as the target. The fine-tuned model then serves as a foundation for the subsequent RL phase.
3.2.2 Phase 2: Reinforcement Learning (RL)
In this phase, we preferentially optimize the fine-tuned model’s performance with RL. As depicted in Figure 2, we train the model to maximize a reward, corresponding to the quality of the generated image, . We define the reward, as a weighted sum of three terms:
Where:
, , and are coefficients denoting the relative importance of , , and (see following points) respectively.
is an aesthetics score obtained from a fully connected neural network trained to take CLIP Radford et al. (2021) image embeddings as input and predict a score representing human aesthetic preferences333https://github.com/christophschuhmann/improved-aesthetic-predictor.
is an alignment score determined by the cosine similarity between CLIP embeddings of the normal prompt and the generated image.
is a fidelity score based on Inception Scores Salimans et al. (2016).
4 Experimental Setup
We use T5444https://hf.co/t5-small Raffel et al. (2020) as our seq2seq language model and Stable Diffusion555https://hf.co/stabilityai/stable-diffusion-2-base Rombach et al. (2022) as our image generator.
4.1 SFT
SFT Dataset Subset
We choose a subset of the Negative Prompts DB dataset, selecting only Stable Diffusion Rombach et al. (2022) posts with 20+ likes. This subset consists of 5,790 samples. The train/validation/test split ratio is 90:5:5.
SFT Training
After hyperparameter search, we select the Adam optimizer Kingma and Ba (2014) with a learning rate of , weight decay, warm-up steps, and an effective batch size of . We fine-tune for epochs with the prefix instruction “generate a negative prompt for:”.
4.2 RL
RL Dataset Subset
We use a more selective subset of 466 samples with 100+ likes from the SFT train and validation splits. We set the new train/validation split ratio to 90:10.
RL Training
After hyperparameter search, we choose the Adam optimizer with a learning rate of and an effective batch size of . We train for epochs and use greedy decoding during evaluation by setting the temperature to . We prioritize aesthetics by setting , , and . We employ Proximal Policy Optimization (PPO) Schulman et al. (2017), a policy gradient algorithm that updates the model’s policy using gradient descent.
4.3 Evaluation
Image Generation Details
We run Stable Diffusion for steps with a guidance scale of on the prompts from the SFT test split. We generate images with 8 different seeds and record the mean of the evaluation metrics.
Evaluation Metrics
We employ these metrics:
Inception Score Salimans et al. (2016): This assesses both the quality and diversity of the generated images.
CLIP Score Hessel et al. (2021): This measures the semantic similarity between images and text, with scores ranging from 0 to 100.
Aesthetics Score666github.com/christophschuhmann/improved-aesthetic-predictor: This gauges the aesthetic appeal of images based on human preferences, with scores ranging from 0 to 10.
Human Evaluation: We have evaluators rank generated images (see Appendix A) then compute the mean rank for each method. Scores range from 1 to the number of methods being evaluated (6, in this case).
Baselines, Previous Methods, & Method Variants
We compare to two baselines: None and Ground Truth, one previous method: Promptist, and three NegOpt variants: SFT-only, RL-only, and SFT+RL. We describe them next:
None does not utilize a negative prompt.
Ground Truth uses the ground-truth negative prompt from the dataset.
Promptist Hao et al. (2022) trains a policy on to augment normal prompts.
SFT-only uses the negative prompt from a model that is trained with SFT but not RL.
RL-only uses the negative prompt from a model trained with RL but not SFT.
SFT+RL combines both SFT and RL.
| Metrics | |||||
| Inception Score (↑) | CLIP Score (↑) | Aesthetics Score (↑) | Mean Human Rank (↓) | ||
| Baselines | None | 5.58 | 32.16 | 6.08 | 3.42 |
| Ground Truth | 6.82 | 30.98 | 6.28 | 2.96 | |
| Promptist | 6.61 | 20.49 | 5.23 | 4.96 | |
| NegOpt | RL-only | 6.98 | 28.37 | 5.76 | 3.94 |
| SFT-only | 7.16 | 30.58 | 6.28 | 2.82 | |
| SFT+RL | 7.08 | 30.88 | 6.30 | 2.90 | |
5 Results and Discussion
5.1 Results
Our results, presented in Table 5, demonstrate the effectiveness of NegOpt compared to the baselines and alternative methods discussed in Section 4.3. In terms of Inception Score, our fidelity measure, we achieve substantial improvements of 24.8% and 8.0% over the None baseline and the top previous method, Promptist, respectively. For the Aesthetics Score, our aesthetics measure, we outperform None and Promptist by 3.5% and 18.6% respectively. Although our CLIP Score, the alignment measure, drops by 4.1% compared to None, this is acceptable as it is our least important concern and the decrease is relatively small. We also observe that RL-only performs worse than SFT-only and to a greater extent SFT+RL.
5.2 Discussion
SFT provides a strong start
The SFT-only approach already demonstrates a significant improvement in image quality, even surpassing NegOpt in Inception Score. This validates the quality of our dataset and the initial part of our approach. However, there is more to consider.
Learning better negative prompts than ground-truth
Interestingly, NegOpt learns better negative prompts than the ground truth in the test set. This is not surprising, as the negative prompts were created by humans, who are essentially performing some manual optimization task. Consequently, NegOpt can leverage patterns in the training data more efficiently than those who produced it. Next, we discuss the importance of RL.
RL optimizes for metrics we care more about
By incorporating RL, we can better control our performance on specific metrics. Based on our settings of , , and from Section 4.2, we prioritized the aesthetics. As a result, SFT+RL outperforms SFT-only in aesthetics. Thus, the addition of RL, and more specifically a multi-component reward function (see Equation 1), enables us to tailor image generation toward what we value most without any substantial degradation in the less valued ones.
6 Conclusion
In this paper, we introduced NegOpt, a novel method for optimizing negative prompts for text-to-image generation tasks. To the best of our knowledge, this is the first approach specifically targeting the optimization of negative prompts. Additionally, we constructed Negative Prompts DB, the first dataset designed for negative prompts. NegOpt demonstrates strong performance in terms of the evaluation metrics: Inception Score, CLIP Score, and Aesthetics Score, showcasing its effectiveness in enhancing image generation. We showed that Supervised Fine-Tuning provides a solid foundation, while further optimization through Reinforcement Learning allows for more targeted improvements in the most relevant metrics. On a broader level, our work highlights the potential of expressing optimization problems in source modalities, such as language, to achieve significant performance gains in target modalities, such as vision. This opens up new avenues for future research in optimizing prompts for non-language tasks.
Acknowledgements
We thank Logan Woudstra for helping out with human evaluation and Lili Mou for providing supplementary computing resources. Finally, we thank Greg Kondrak for general guidance in our research.
Ethics Statement
In developing our NegOpt method for optimizing negative prompts in image generation, we are committed to upholding ethical principles and ensuring responsible use. We acknowledge potential ethical concerns, including dataset bias, potential misuse, privacy concerns, ethical reinforcement learning, and transparency and explainability. To address these concerns, we strive to:
1.
Assess and mitigate biases in the data, promoting fairness and inclusivity.
2.
Encourage responsible use by providing guidelines and recommendations.
3.
Respect user privacy by adhering to data protection regulations.
4.
Explore incorporating ethical considerations into the reinforcement learning reward function.
5.
Improve the transparency and explainability of our method.
We believe that addressing these ethical concerns contributes to the development of ethically sound and beneficial image-generation technology. We encourage future research to consider these ethical aspects and collaborate in fostering responsible innovation in the field.
Limitations
Dataset bias
The Negative Prompts DB dataset may contain inherent biases, which could lead to biased image generation outcomes.
Potential misuse
Our method could be misused to generate harmful content or misinformation, emphasizing the need for guidelines and safeguards.
Privacy concerns
Although we do not use personally identifiable information, there may still be privacy concerns related to using user-generated content from Playground AI.
Ethical reinforcement learning
The reward function in our method may not fully capture ethical implications, necessitating the exploration of incorporating ethical considerations.
Transparency and explainability
The deep learning models used in our method may raise concerns about the interpretability of the generated negative prompts and resulting images.
Appendix
Appendix A Human Evaluation
The instruction provided to human evaluators is: “For each sample, you have six images and a prompt. You are to rank all images by two criteria of equal importance: (1) Beauty/aesthetic pleasantness, and (2) Relevance to the prompt. The ranking should be an ordered list of numbers (eg. 1,5,6,4,2,3; if you think the first image is the best and the third image is the worst).”
Appendix B NegOpt Examples
See Table 3 for NegOpt vs. all baselines and methods.
Prompt |
Negative Prompt |
Image |
lushill style, Woman head and shoulders portrait. style by Sir John Tenniel, Beatrix Potter, Brian Froud, Enki Bilal, Pascal Campion. Lovely flower ornaments on the wall behind her. |
None |
Ground Truth |
open-mouth, hands, duplicates, blur haze |
Promtist |
RL-only |
, Brian Froud, Enki Bilal, Pascal Campion. |
SFT-only |
elongated body, cropped image, red eyes, out of frame, draft, deformed hands, signatures, big hair, twisted fingers, double image, long neck, malformed hands, multiple heads, extra limb, ugly… |
SFT+RL |
elongated body, cropped image, red eyes, out of frame, draft, deformed hands, signatures, big hair, twisted fingers, double image, long neck, malformed hands, multiple heads, extra limb, ugly… |
elven architecture a small cottage intricate detailed wonder ancient woods background spring flowers vines, beautiful garden with brook in background, golden sunlight and sunbeams photographic render, smooth sharp focus… |
None |
Ground Truth |
ugly, tiling, poorly drawn hands, poorly drawn feet, poorly drawn face, out of frame, extra limbs, disfigured, deformed, body out of frame, blurry, bad anatomy, blurred, watermark, grainy, signature, cut off, draft |
Promtist |
RL-only |
, artillery, sparks, racks, motherboard, by pascal blanche rutkowski,, cinematic, 4k, epic Steven Spielberg movie still, sharp focus, emitting diodes, smoke, artillery, sparks, racks, system unit, motherboard… |
SFT-only |
elongated body, cropped image, red eyes, out of frame, draft, deformed hands, signatures, big hair, twisted fingers, double image, long neck, malformed hands, multiple heads, extra limb, ugly, poorly drawn hands… |
SFT+RL |
elongated body, cropped image, red eyes, out of frame, draft, deformed hands, signatures, big hair, twisted fingers, double image, long neck, malformed hands, multiple heads, extra limb, ugly, poorly drawn hands… |
![[Uncaptioned image]](extracted/5465360/assets/girl.jpg)
![[Uncaptioned image]](extracted/5465360/assets/girl_neg_true.jpg)
![[Uncaptioned image]](extracted/5465360/assets/girl_neg_pred.jpg)